OpenStudy (anonymous):

Is anyone able to help with showing the method for this? sin 2x - sin x = 0

OpenStudy (anonymous):

2sinx cosx-sinx=0

OpenStudy (unklerhaukus):

\[\sin 2x - \sin x = 0\] \[\sin 2x = \sin x \] \[f(x)=\sin x\qquad\qquad g(x)=\sin 2x\]